# HOPSE: Scalable Higher-Order Positional and Structural Encoder for Combinatorial Representations

### v1

@@ 5.2 Experimental Results
Table 1: Accuracy on benchmark datasets. Results are shown as mean $\pm$ standard deviation. The best results are highlighted in bold, and those within one standard deviation of the best are in blue.
@@
Graph Tasks Topological Tasks Model MUTAG ($\uparrow$) PROTEINS ($\uparrow$) NCI1 ($\uparrow$) NCI109 ($\uparrow$) ZINC ($\downarrow$) NAME ($\uparrow$) ORIENT ($\uparrow$) $\beta_{1}$ ($\uparrow$) $\beta_{2}$($\uparrow$) Graph GCN 69.79 $\pm$ 6.80 75.70 $\pm$ 2.14 72.86 $\pm$ 0.69 72.20 $\pm$ 1.22 0.62 $\pm$ 0.01 42.14 $\pm$ 2.72 47.94 $\pm$ 0.00 46.86 $\pm$ 4.50 0.00 $\pm$ 0.00 GAT 72.77 $\pm$ 2.77 76.34 $\pm$ 1.66 75.00 $\pm$ 0.99 73.80 $\pm$ 0.73 0.61 $\pm$ 0.01 18.09 $\pm$ 0.65 47.94 $\pm$ 0.00 7.45 $\pm$ 0.05 0.00 $\pm$ 0.00 GIN 79.57 $\pm$ 6.13 75.20 $\pm$ 3.30 74.26 $\pm$ 0.96 74.42 $\pm$ 0.70 0.57 $\pm$ 0.04 76.14 $\pm$ 0.14 56.28 $\pm$ 0.45 88.13 $\pm$ 0.00 0.93 $\pm$ 1.21 Simplicial SCN 73.62 $\pm$ 6.13 75.27 $\pm$ 2.14 74.49 $\pm$ 1.03 75.70 $\pm$ 1.04 0.53 $\pm$ 0.04 79.48 $\pm$ 1.36 69.55 $\pm$ 0.97 76.45 $\pm$ 3.06 5.45 $\pm$ 2.31 SCCNN 76.17 $\pm$ 6.63 74.19 $\pm$ 2.86 76.60 $\pm$ 1.75 77.12 $\pm$ 1.07 0.36 $\pm$ 0.02 95.08 $\pm$ 0.56 86.29 $\pm$ 1.23 90.20 $\pm$ 0.20 65.82 $\pm$ 2.70 SaNN 79.57 $\pm$ 3.56 75.05 $\pm$ 3.52 68.75 $\pm$ 1.64 68.17 $\pm$ 0.88 0.55 $\pm$ 0.01 81.76 $\pm$ 1.37 61.65 $\pm$ 0.55 88.46 $\pm$ 0.09 39.22 $\pm$ 2.80 GCCN 85.96 $\pm$ 4.66 74.91 $\pm$ 2.51 76.85 $\pm$ 1.50 75.76 $\pm$ 1.28 0.35 $\pm$ 0.01 86.76 $\pm$ 1.27 76.60 $\pm$ 1.67 84.20 $\pm$ 4.80 41.82 $\pm$ 20.19 HOPSE-M 86.38 $\pm$ 6.31 75.13 $\pm$ 3.66 77.35 $\pm$ 0.81 76.34 $\pm$ 1.40 0.39 $\pm$ 0.02 91.50 $\pm$ 1.45 80.68 $\pm$ 1.72 90.26 $\pm$ 0.55 71.69 $\pm$ 1.50 HOPSE-G 85.53 $\pm$ 2.77 76.27 $\pm$ 2.10 77.70 $\pm$ 1.40 76.36 $\pm$ 1.75 0.43 $\pm$ 0.01 81.75 $\pm$ 1.26 62.17 $\pm$ 0.98 88.28 $\pm$ 0.08 35.37 $\pm$ 2.25 Cell CCCN 77.02 $\pm$ 9.32 73.33 $\pm$ 2.30 76.67 $\pm$ 1.48 75.35 $\pm$ 1.50 0.34 $\pm$ 0.02 - - - - CWN 80.43 $\pm$ 1.78 76.13 $\pm$ 2.70 73.93 $\pm$ 1.87 73.80 $\pm$ 2.06 0.34 $\pm$ 0.01 - - - - GCCN 86.38 $\pm$ 6.49 74.41 $\pm$ 1.77 78.23 $\pm$ 1.47 77.19 $\pm$ 0.21 0.19 $\pm$ 0.00 - - - - HOPSE-M 88.09 $\pm$ 5.55 76.99 $\pm$ 2.63 78.09 $\pm$ 1.44 77.27 $\pm$ 1.01 0.25 $\pm$ 0.01 - - - - HOPSE-G 88.09 $\pm$ 1.90 76.42 $\pm$ 2.19 79.07 $\pm$ 0.65 77.81 $\pm$ 1.26 0.21 $\pm$ 0.01 - - - -
Accuracy comparison.
Table 1 presents a performance comparison across topological domains: graph, simplicial, and cell. The proposed HOPSE-M and HOPSE-G models demonstrate consistently strong performance, particularly in the simplicial domain, answering both Q1. and Q2. affirmatively. On standard graph classification benchmarks used in TDL, both HOPSE variants match or outperform most baselines, with HOPSE-M ranking among the top models and achieving state-of-the-art results in several cases (PROTEINS, NCI1, NCI109). While the performance gap on ZINC highlights the advantage of message-passing in fine-grained molecular regression tasks, HOPSE still remains competitive without relying on the HOMP procedure. The most significant gains are observed on the pure topological datasets. HOPSE-M delivers the strongest results on $\beta_{1}$ and $\beta_{2}$, outperforming all baselines, including models with explicit simplicial message passing. This suggests that HOPSE can effectively encode and leverage topological structure. In summary, the HOPSE framework achieves strong performance across both graph and synthetic topological benchmarks, demonstrating that a carefully designed structural encoding pipeline can approximate or surpass topological message-passing methods. Detailed analysis of the results is provided in Appendix F.
Empirical expressivity.
Table 2: Accuracy on CSL ($\uparrow$).
Model Mean Max Min MP-GNNs[9] 10.00 $\pm$ 0.00 10.00 10.00 SaNN 26.00 $\pm$ 6.83 33.33 20.00 SCCNN 25.33 $\pm$ 7.67 33.33 16.67 GCCN 52.78 $\pm$ 34.28 100.00 3.33 3WLGNN [32] 97.80 $\pm$ 10.92 100.00 30.00 HOPSE-M 100.00 $\pm$ 0.00 100.00 100.00 HOPSE-G 98.00 $\pm$ 4.47 100.00 90.00
Table 2 reports the mean, minimum, and maximum test classification accuracy over five folds for the best-performing baselines, covering all domains. The results show that SaNN and SCCNN slightly outperform baseline MP-GNNs, while GCCN performs better on the cell domain but still falls short of 3WLGNN. HOPSE-M perfectly distinguishes all isomorphism classes, and HOPSE-G achieves the second-best position ( Q3.).
Computation time comparison.
To highlight the efficiency of HOPSE, Table 3 reports the total runtime in seconds. Notably, either HOPSE-G or HOPSE-M achieves the fastest overall runtime across both simplicial and cell domains. On graph datasets, the speedup can reach up to 7x—for example, HOPSE-G runs in 130 seconds compared to 904 seconds for GCCN on NCI1, while also delivering better performance. On average, HOPSE-M and HOPSE-G outperform the most competitive baselines with speedups of 1.5x and 3x over SCCN, and 3x and 5x over GCCN, respectively, while achieving top results on 6 out of 9 datasets. Overall, both models significantly reduce runtime while delivering competitive or superior accuracy (Q4.).
Table 3: Runtime in seconds. Results are shown as mean $\pm$ standard deviation. The fastest model per domain is in bold and those that are within one standard deviation of the best are in blue.
@@
Graph Tasks Topological Tasks Model MUTAG PROTEINS NCI1 NCI109 ZINC NAME ($\uparrow$) ORIENT $(\beta_{0},\beta_{1},\beta_{2})$ Graph GCN 4 $\pm$ 1 8 $\pm$ 2 53 $\pm$ 20 37 $\pm$ 9 147 $\pm$ 28 3538 $\pm$ 38 3534 $\pm$ 49 7250 $\pm$ 234 GAT 4 $\pm$ 1 8 $\pm$ 2 57 $\pm$ 17 56 $\pm$ 9 171 $\pm$ 65 3589 $\pm$ 84 3544 $\pm$ 44 3612 $\pm$ 52 GIN 5 $\pm$ 1 9 $\pm$ 2 61 $\pm$ 24 50 $\pm$ 8 168 $\pm$ 107 3568 $\pm$ 44 3536 $\pm$ 46 3674 $\pm$ 176 Simplicial SCN 8 $\pm$ 2 43 $\pm$ 13 276 $\pm$ 63 226 $\pm$ 66 1210 $\pm$ 572 5115 $\pm$ 653 4773 $\pm$ 193 6992 $\pm$ 2049 SCCNN 14 $\pm$ 3 54 $\pm$ 11 307 $\pm$ 83 354 $\pm$ 106 2060 $\pm$ 408 5521 $\pm$ 319 6920 $\pm$ 1524 6720 $\pm$ 210 SaNN 9 $\pm$ 2 50 $\pm$ 12 250 $\pm$ 18 211 $\pm$ 66 1263 $\pm$ 252 5572 $\pm$ 1053 4619 $\pm$ 103 5220 $\pm$ 421 GCCN 61 $\pm$ 7 66 $\pm$ 21 904 $\pm$ 180 538 $\pm$ 39 3603 $\pm$ 475 5898 $\pm$ 1961 4564 $\pm$ 631 7689 $\pm$ 2522 HOPSE-M 7 $\pm$ 2 38 $\pm$ 8 164 $\pm$ 79 219 $\pm$ 26 1061 $\pm$ 405 3596 $\pm$ 519 3409 $\pm$ 162 5377 $\pm$ 185 HOPSE-G 5 $\pm$ 0 33 $\pm$ 8 130 $\pm$ 22 107 $\pm$ 28 478 $\pm$ 107 3324 $\pm$ 1113 2795 $\pm$ 112 3876 $\pm$ 1141 Cell CCCN 12 $\pm$ 2 42 $\pm$ 7 372 $\pm$ 109 272 $\pm$ 21 1621 $\pm$ 142 - - - CWN 11 $\pm$ 1 54 $\pm$ 18 302 $\pm$ 63 295 $\pm$ 46 1390 $\pm$ 97 - - - GCCN 61 $\pm$ 18 59 $\pm$ 18 523 $\pm$ 119 386 $\pm$ 76 3301 $\pm$ 440 - - - HOPSE-M 7 $\pm$ 1 37 $\pm$ 7 147 $\pm$ 24 122 $\pm$ 18 1313 $\pm$ 132 - - - HOPSE-G 7 $\pm$ 1 32 $\pm$ 5 153 $\pm$ 8 104 $\pm$ 38 1230 $\pm$ 527 - - -
Choice of ${\mathcal{N}}_{C}$.
The neighborhood set ${\mathcal{N}}_{C}$ is a critical design choice in the HOPSE framework, influencing both the expressivity and computational behavior of the model. Performed ablation studies (Appendix F) reveal that richer (i.e. larger and diverse) neighborhood configurations consistently yield top or near-top accuracy across complex datasets—e.g. MANTRA’s topological tasks, ZINC—as shown in Tables 6 and 7. Importantly, these gains come without significant runtime penalties: as demonstrated in Tables 10 and 11, per-epoch runtime remains almost stable, and total runtime often decreases due to faster convergence. Notably, the cell domain strongly aligns with expressivity and efficiency, especially on molecular tasks. These findings suggest a practical guideline: practitioners aiming for strong performance without exhaustive search over all neighborhood configurations can safely opt for larger ${\mathcal{N}}_{C}$ sets. This choice is especially advantageous for complex tasks, balancing predictive power, convergence speed, and runtime efficiency (Q1., Q2., Q4.).
@@ Conclusions.
This work introduces HOPSE, a scalable framework designed to extract and leverage higher-order structural and positional information for arbitrary combinatorial representations. Theoretically, HOPSE matches the expressivity of state-of-the-art TNNs while avoiding the HOMP paradigm, thereby substantially reducing the computational cost of training and inference. Moreover, model scaling is linear with dataset size. These claims are supported by experiments on molecular graph benchmarks, as well as expressivity and topological tasks, where both instantiations—HOPSE-M and HOPSE-G—match or surpass state-of-the-art performance while achieving up to 7×7$\times$7 × speed-ups over HOMP-based models. By offering a compelling trade-off between scalability and expressivity, this work not only enables broader applicability of TDL but also facilitates the integration of higher-order interactions into other areas of deep learning—mirroring the widespread adoption of graph-based methods [47, 44, 24, 25].
@@
A deeper investigation of HOPSE encodings—along with a comprehensive comparison against alternative methods across diverse domains—lies beyond the scope of this work but remains an important and intriguing direction for future research. Other promising avenues include exploring graph-based strategies for generating higher-order positional and structural encodings, or combining encodings derived from multiple topological domains for enhanced expressivity. Additionally, this work does not address whether integrating the proposed encodings into GNNs, TNNs, or transformer-based architectures leads to substantial improvements across a range of downstream tasks. Finally, and from a more practical standpoint, HOPSE-G realization of the framework leverages a pretrained model to extract latent structural and positional features, but it would be interesting to explore if a joint end-to-end training may further enhance performance –even at the cost of additional computational overhead.
@@ A.1 Topological domains
A topological domain is a graph generalization that captures pairwise and higher-order relationships between entities [7, 4]. When working with topological domains, two key properties come into play: set-type relations and hierarchical structures represented by rank functions [23, 38].
@@
Simplicial complexes extend graphs by incorporating hierarchical part-whole relationships through the multi-scale construction of cells. In this structure, nodes correspond to rank 00-cells, which can be combined to form edges (rank $1$-cells). Edges can then be grouped to form faces (rank 2 cells), and faces can be combined to create volumes (rank $3$-cells), continuing in this manner. Consequently, the faces of a simplicial complex are triangles, volumes are tetrahedra, and higher-dimensional cells follow the same pattern. A key feature of simplicial complexes is their strict hierarchical structure, where each $k$-dimensional simplex is composed of $(k-1)$-dimensional simplices, reinforcing a strong sense of hierarchy across all levels.
@@
Cell complexes provide a hierarchical interior-to-boundary structure, offering clear topological and geometric interpretations, but they are not based on set-type relations. Unlike simplicial complexes, cell complexes are not limited to simplexes; faces can involve more than three nodes, allowing for a more flexible representation. This increased flexibility grants cell complexes greater expressivity compared to simplicial complexes [10, 8].
@@
A featured graph is a graph whose nodes or edges are equipped with feature functions [41]. Similarly, each topological element (e.g., simplex or cell) can carry feature vectors to represent some information.

### v2

@@ 5.2 Experimental Results
Table 1: Performance (ACC,MAE) and Total training time in seconds (T) on Graph Tasks.
@@
Model Dataset MUTAG PROTEINS NCI1 NCI109 ZINC-12K ACC ($\uparrow$) T ($\downarrow$) ACC ($\uparrow$) T ($\downarrow$) ACC ($\uparrow$) T ($\downarrow$) ACC ($\uparrow$) T ($\downarrow$) MAE ($\downarrow$) T ($\downarrow$) Graph GCN $69.79\pm 0.06$.80) $4\pm 1.$ $75.70\pm 0.02$.14) $8\pm 2.$ $72.86\pm 0.00$.69) $53\pm 20.$ $72.20\pm 0.01$.22) $37\pm 9.$ $0.62\pm 0.00$.01) $146\pm 28.$ GCN+PSEs $84.26\pm 0.03$.56) $3\pm 1.$ $75.77\pm 0.02$.98) $23\pm 5.$ $74.88\pm 0.00$.54) $106\pm 47.$ $74.70\pm 0.00$.79) $122\pm 39.$ $0.43\pm 0.00$.03) $903\pm 40.$ GAT $2.77\pm 0.02.77$) $4\pm 1.$ $76.34\pm 0.01$.66) $8\pm 2.$ $75.00\pm 0.01$.00) $57\pm 17.$ $73.80\pm 0.00$.73) $56\pm 9.$ $0.61\pm 0.00$.01) $171\pm 64.$ GAT+PSEs $80.43\pm 0.05$.08) $5\pm 1.$ $75.48\pm 0.03$.44) $36\pm 15.$ $77.86\pm 0.01$.02) $210\pm 28.$ $75.57\pm 0.00$.60) $212\pm 89.$ $0.39\pm 0.00$.02) $708\pm 83.$ Simplicial SCCNN $76.17\pm 0.06$.63) $14\pm 3.$ $74.19\pm 0.02$.86) $53\pm 11.$ $76.60\pm 0.01$.75) $307\pm 83.$ $77.12\pm 0.01$.07) $354\pm 106.$ $0.36\pm 0.00$.02) $2060\pm 408.$ GCCN $85.96\pm 0.04$.66) $61\pm 7.$ $74.91\pm 0.02$.51) $66\pm 21.$ $76.85\pm 0.01$.50) $904\pm 180.$ $75.76\pm 0.01$.28) $538\pm 39.$ $0.35\pm 0.00$.01) $3603\pm 475.$ HOPSE-M (Our) $86.38\pm 0.06$.31) $10\pm 3.$ $75.13\pm 0.03$.66) $44\pm 17.$ $77.35\pm 0.00$.81) $201\pm 88.$ $76.34\pm 0.01$.40) $179\pm 79.$ $0.39\pm 0.00$.02) $1332\pm 721.$ HOPSE-G (Our) $85.53\pm 0.02$.77) $5\pm 1.$ $76.27\pm 0.02$.10) $33\pm 8.$ $77.70\pm 0.01$.40) $130\pm 22.$ $76.36\pm 0.01$.75) $107\pm 28.$ $0.43\pm 0.00$.01) $478\pm 107.$ Cell GCCN $86.38\pm 0.06$.49) $61\pm 18.$ $4.41\pm 0.01.77$) $59\pm 18.$ $78.23\pm 0.01$.47) $523\pm 119.$ $77.19\pm 0.00$.21) $386\pm 76.$ $0.19\pm 0.00$.01) $3301\pm 440.$ HOPSE-M (Our) $88.09\pm 0.05$.55) $7\pm 3.$ $76.99\pm 0.02$.63) $35\pm 17.$ $78.09\pm 0.01$.44) $156\pm 90.$ $77.27\pm 0.01$.01) $147\pm 81.$ $0.24\pm 0.00$.01) $760\pm 533.$ HOPSE-G (Our) $88.09\pm 0.01$.90) $7\pm 1.$ $76.42\pm 0.02$.19) $32\pm 5.$ $79.07\pm 0.00$.65) $153\pm 8.$ $77.81\pm 0.01$.26) $104\pm 38.$ $0.21\pm 0.00$.01) $1230\pm 527.$
Table 2: Performance (F1) and Total training time in seconds (T) on Topological Tasks.
@@
Model Task NAME ORIENT $\beta_{1}$ $\beta_{2}$ F1 ($\uparrow$) T ($\downarrow$) F1 ($\uparrow$) T ($\downarrow$) F1 ($\uparrow$) T ($\downarrow$) F1 ($\uparrow$) T ($\downarrow$) Graph GIN $76.14\pm 0.00$.14) $3568\pm 44.$ $56.28\pm 0.00$.45) $3536\pm 46.$ $88.13\pm 0.00$.00) $3674\pm 176.$ $0.93\pm 0.01$.21) $3674\pm 176.$ GIN+PSEs $94.05\pm 0.00$.15) $2174\pm 306.$ $82.53\pm 0.00$.78) $2128\pm 247.$ $90.28\pm 0.00$.05) $3011\pm 562.$ $68.43\pm 0.00$.59) $3011\pm 562.$ Simplicial SCCNN $95.08\pm 0.00$.56) $5521\pm 319.$ $86.29\pm 0.01$.23) $3536\pm 46.$ $90.20\pm 0.00$.20) $6720\pm 10.$ $65.82\pm 0.02$.70) $6720\pm 10.$ GCCN $86.76\pm 0.01$.27) $5898\pm 1961.$ $76.60\pm 0.01$.67) $4564\pm 631.$ $84.20\pm 0.04$.80) $7689\pm 2522.$ $41.82\pm 0.20$.19) $7689\pm 2522.$ HOPSE-M (Our) $96.27\pm 0.00$.28) $1530\pm 1366.$ $89.24\pm 0.00$.22) $1886\pm 1449.$ $91.02\pm 0.01$.43) $3179\pm 1756.$ $82.29\pm 0.00$.46) $3179\pm 1756.$ HOPSE-G (Our) $81.75\pm 0.01$.26) $3324\pm 1113.$ $62.17\pm 0.00$.98) $2795\pm 112.$ $88.28\pm 0.00$.08) $3876\pm 1141.$ $35.37\pm 0.02$.25) $3876\pm 1141.$
HOMP is not a hard requirement for performance.
Table˜1 and Table˜2 present the performance in graph and topological tasks respectively. The performance metric is the left most column for each dataset. In both scenarions, HOPSE-M and HOPSE-G models demonstrate consistently strong performance. This results illustrates the both Q1. and Q2. are negative, HOMP is not a requirement to capture higher-order relationships in traditional TDL benchmarks or in tasks related to topological properties. On standard graph classification benchmarks used in TDL, both HOPSE variants match or outperform most baselines, with HOPSE-M ranking among the top models and achieving state-of-the-art results in several cases (PROTEINS, NCI1, NCI109). While the performance gap on ZINC highlights the advantage of message-passing in fine-grained molecular regression tasks, HOPSE still remains competitive without relying embbedings generated with HOMP. The most significant gains are observed on the pure topological datasets. HOPSE-M delivers the strongest results on $\beta_{1}$ and $\beta_{2}$, outperforming all baselines, including models with explicit simplicial message passing. This suggests that HOPSE can effectively encode and leverage topological structure. It is also noteworthy that Table˜2 shows that GIN+PSEs always outperforms GCCNs and in $\beta_{1},\beta_{2}$ it outperforms SCCN. This results put into question if HOMP is usefull at all for topological tasks. For performance over all models refer to Appendix˜G, for ablations over different neighborhoods to Appendix˜H and for ablations over PSEs to Appendix˜J.
Speed does not need to compromise performance.
Tables˜2 and 1 show the total runtime of training in the rightmost column of each dataset.  [^5] This does not included preprocessing times, see Section I.3 for more in depth explanation.. In the category of models operating over simplicial or cell complexes, a HOPSE realization always achieves the fastest overall runtime. On graph datasets, the speedup can reach up to $7\times$—for example, HOPSE-G runs in 130 seconds compared to 904 seconds for GCCN on NCI1, while also delivering better performance. HOPSE’s efficiency becomes apparant as tasks scale in size and complexity. In NCI1, NCI109 and ZINC it is from $3\times$ to $7\times$ faster than other models with comparable or better performance. In topological tasks this is exacerbated, where HOPSE-M is up to $1.5\times$ more efficienct than GIN+PSEs while outperforming all other models. These results show that the theoretical complexity does result in empirically reduced training time, even when comparing to the case of GNNs. For runtimes over all models refer to Appendix˜G, for ablations over different neighborhoods to Appendix˜I and for model sizes to Appendix˜K.
@@ Conclusions.
This work introduces HOPSE, a scalable and compute efficient framework designed to extract and leverage structural information of the Hasse decomposition to characterize arbitrary combinatorial representations. HOPSE performs competitively with HOMP without performing any type of message passing. Moreover, model scaling is linear with dataset size due parallel learning of structure and attributes. These claims are supported by experiments on molecular graph benchmarks, as well as and topological tasks, where both instantiations—HOPSE-M and HOPSE-G—match or surpass state-of-the-art performance while achieving up to $7\times$ speed-ups over HOMP-based models. By providing a characterization of a complex object via it’s constituting objects, Hasse graphs, HOPSE provides a new method for solving topological tasks. More importantly, in puts into question the necessity of HOMP. The cost of operating over augmented neighborhood structures blows up complexity without improving the model’s ability to learn the topology of the data. In a novel comparison, the ablations over GNNs with PSEs agains HOMP show that in several cases there is no need for the latter.
Nevertheless, this work not only enables broader applicability of TDL but also facilitates the integration of higher-order interactions into other areas of deep learning—mirroring the widespread adoption of graph-based methods (Zhou et al., 2023; Wang et al., 2025; Han et al., 2022; 2023).
@@
A deeper investigation of HOPSE encodings—along with a comprehensive comparison against alternative methods across diverse domains—lies beyond the scope of this work but remains an important and intriguing direction for future research. Other promising avenues include exploring graph-based strategies for generating higher-order positional and structural encodings, or better methods for characterizing Hasse graphs like the ones based on homomorphism counts. Additionally, this work does not address whether directly manufacturing encodings on a combinatorial complex provides a better characterization the joint characterization using it’s Hasse decomposition.Finally, and from a more practical standpoint, HOPSE-G realization of the framework leverages a pretrained model to extract latent structural and positional features, but it would be interesting to explore if a joint end-to-end training may further enhance performance –even at the cost of additional computational overhead.
@@ A.1 Topological domains
A topological domain is a graph generalization that captures pairwise and higher-order relationships between entities [Bick et al., 2023, Battiston et al., 2021]. When working with topological domains, two key properties come into play: set-type relations and hierarchical structures represented by rank functions [Hajij et al., 2023, Papillon et al., 2023].
@@
Simplicial complexes extend graphs by incorporating hierarchical part-whole relationships through the multi-scale construction of cells. In this structure, nodes correspond to rank 0-cells, which can be combined to form edges (rank $1$-cells). Edges can then be grouped to form faces (rank 2 cells), and faces can be combined to create volumes (rank $3$-cells), continuing in this manner. Consequently, the faces of a simplicial complex are triangles, volumes are tetrahedra, and higher-dimensional cells follow the same pattern. A key feature of simplicial complexes is their strict hierarchical structure, where each $k$-dimensional simplex is composed of (k1)(k-1)-dimensional simplices, reinforcing a strong sense of hierarchy across all levels.
@@
Cell complexes provide a hierarchical interior-to-boundary structure, offering clear topological and geometric interpretations, but they are not based on set-type relations. Unlike simplicial complexes, cell complexes are not limited to simplexes; faces can involve more than three nodes, allowing for a more flexible representation. This increased flexibility grants cell complexes greater expressivity compared to simplicial complexes Bodnar et al. [2021], Bodnar [2023].
@@
A featured graph is a graph whose nodes or edges are equipped with feature functions [Sanchez-Lengeling et al., 2021]. Similarly, each topological element (e.g., simplex or cell) can carry feature vectors to represent some information.